# Beyond Ensemble Averages: Leveraging Climate Model Ensembles for Subseasonal Forecasting

### v3

@@ Abstract
Producing high-quality forecasts of key climate variables, such as temperature and precipitation, on subseasonal time scales has long been a gap in operational forecasting. This study explores an application of machine learning (ML) models as post-processing tools for subseasonal forecasting. Lagged numerical ensemble forecasts (i.e., an ensemble where the members have different initial dates) and observational data, including relative humidity, pressure at sea level, and geopotential height, are incorporated into various ML methods to predict monthly average precipitation and two-meter temperature two weeks in advance for the continental United States. Regression, quantile regression, and tercile classification tasks using linear models, random forests, convolutional neural networks, and stacked models (a multi-model approach based on the prediction of the individual ML models) are considered. Unlike previous ML approaches that often use ensemble mean alone, we leverage information embedded in the ensemble forecasts to enhance prediction accuracy. Additionally, we investigate extreme event predictions that are crucial for planning and mitigation efforts. Considering ensemble members as a collection of spatial forecasts, we explore different approaches to address spatial variability. Trade-offs between different approaches may be mitigated with model stacking. Our proposed models outperform standard baselines such as climatological forecasts and ensemble means. This paper further includes an investigation of feature importance, trade-offs between using the full ensemble or only the ensemble mean, and different modes of accounting for spatial variability.
@@ 1 Introduction
The National Centers for Environmental Prediction (NCEP), part of the National Oceanic and Atmospheric Administration (NOAA), currently issues a “week 3-4 outlook” for the contiguous United States (CONUS). [^1] https://www.cpc.ncep.noaa.gov/products/predictions/WK34/ The NCEP outlooks are constructed using a combination of dynamical and statistical forecasts, with statistical forecasts based largely on how conditions in the past have varied (linearly) with indices of the El Niño-Southern Oscillation (ENSO), Madden-Julian Oscillation (MJO), and global warming (i.e., the 30-year trend). There exists great potential to advance subseasonal forecasting (SSF) using machine learning (ML) techniques. Haupt et al. [10] provides an overview of ML methods as post-processing of numerical weather predictions. Vannitsem et al. [11] highlight the crucial role of statistical post-processing techniques, including ML methods, in national meteorological services. They discuss theoretical developments and operational applications, current challenges, and potential future directions, with a particular focus on translating of research findings into operational practices. A real-time forecasting competition called the Subseasonal Climate Forecast Rodeo [12], sponsored by the Bureau of Reclamation in partnership with NOAA, USGS, and the U.S. Army Corps of Engineers, illustrated that teams using ML techniques can outperform forecasts from NOAA’s operational seasonal forecast system.
Here we present work focused on developing ML-based forecasts that leverage lagged ensembles (an ensemble whose members are initialized from a succession of different start dates) of forecasts produced by NCEP in addition to observed data and other features. Previous studies, including successful methods in the Rodeo competition (e.g., Hwang et al. [13]), incorporated the ensemble mean as a feature in their ML systems, but did not use any other information about the ensemble. In other words, variations among the ensemble members are not reflected in the training data or incorporated into the learned model. In contrast, this paper demonstrates that the full ensemble contains important information for subseasonal forecasting outside the ensemble mean. Specifically, we consider the test case of predicting monthly 2-meter temperatures and precipitation two weeks in advance over 3000 locations over the continental United States using physics-based predictions, such as NCEP-CFSv2 hindcasts [14, 15], using an ensemble of 24 distinct forecasts. We repeat this experiment for the Global Modeling and Assimilation Office from the National Aeronautics and Space Administration (NASA-GMAO) ensemble, which has 11 ensemble members [16].
@@
  • We train a variety of ML models (including neural networks, random forests, linear regression, and model stacking) that input all ensemble member predictions as features in addition to observations of geopotential heights, relative humidity, precipitation, and temperature from past months to produce new forecasts with higher accuracy than the ensemble mean; forecast accuracy is measured with a variety of metrics (Section 7). These models are considered in the context of regression, quantile regression, and tercile classification. Systematic experiments are used to characterize the influence of individual ensemble members on predictive skill (Section 8).
  • The collection of ML models employed allows us to consider different modes of accounting for spatial variability. ML models can account for spatial correlations among both features and targets; as an example, when predicting Chicago precipitation, our models can leverage not only information about Chicago, but also about neighboring regions. Specifically, we consider the following learning frameworks: (a) training a predictive model for each spatial location independently; (b) training a predictive model that inputs the spatial location as a feature and hence can be applied to any single spatial location; (c) training a predictive model for the full spatial map of temperature or precipitation – i.e., predicting an outcome for all spatial locations simultaneously. ML models present various options for accounting for spatial variability, each with distinct advantages and disadvantages. Our application of model stacking (an ML technique where multiple models are combined, with their predictions used as input features for another model that produces the final prediction) allows our final learned model to exploit the advantages of each method.
@@
This paper is organized as follows: Section 2 discusses related work, Section 3 introduces data used in the experiments, Section 4 describes forecasting problems, while baselines and learning-based methods are described in Section 5, experimental setup and evaluation metrics are given in Section 6. Finally, we present our results in Section 7, and discuss them in Section 8. Conclusions and directions for future work are given in Section 9.
@@ 2 Related work
While statistical models were common for weather prediction in the early days of weather forecasting [17], forecasts using physics-based dynamic system models have been carried out since the 1980s and have been the dominant forecasting method in climate prediction centers since the 1990s [6]. Many physics-based forecast models are used both in academic research and operationally. Such systems often produce ensembles of forecasts – e.g., the result of running a physics-based simulation multiple times with different initial conditions or different parameters, and are a mainstay of operational forecast centers around the globe.
@@
Some ML algorithms for subseasonal forecasting use purely observational data (i.e., not using any physics-based ensemble forecasts). He et al. [35] focuses on analyzing different ML methods, including Gradient Boosting trees and Deep Learning (DL) for SSF. They propose a careful construction of feature representations of observational data and show that ML methods are able to outperform a climatology baseline, i.e., predictions corresponding to the 30-year average at a given location and time. This conclusion is made based on comparing the relative $R^{2}$ scores for the ML approaches and climatology. Srinivasan et al. [36] proposes a Bayesian regression model that exploits spatial smoothness in the data.
Other works use the ensemble mean as a feature in their ML models. For example, in the subseasonal forecasting Rodeo [12], a prediction challenge for temperature and precipitation at weeks 3-4 and 5-6 in the western U.S. sponsored by NOAA and the U.S. Bureau of Reclamation, simple yet thoughtful statistical models consistently outperform NOAA’s dynamical systems forecasts. In particular, the winning approach uses a stacked model from two nonlinear regression models, a selection of climate variables such as temperature, precipitation, sea surface temperature, sea ice concentration, and a collection of physics-based forecast models including the ensemble mean from various modeling centers in NMME. From the local linear regression with multitask feature selection model analysis, the ensemble mean is the first- or second-most important feature for forecasting, especially for precipitation. He et al. [37] perform a comparison of modern ML models that use data from the Subseasonal Experiment (SubX) project for SSF in the western contiguous United States. The experiments show that incorporating the ensemble mean as an input feature to ML models leads to a significant improvement in forecasting performance, but that work does not explore the potential value of individual ensemble members aside from the ensemble mean. Grönquist et al. [38] notes that physics-based ensembles are computationally demanding to produce, and proposes an ML method can input a subset of ensemble forecasts and generate an estimate of the full ensemble; they observe that the output ensemble estimate has more prediction skill than the original ensemble. Loken et al. [39] analyze the forecast skill of random forests leveraging the ensemble members for next-day severe weather prediction compared to only using the ensemble mean. However, their results only cover forecasts with a lead time of up to 48 hours, so it is unclear if their methods would have had success in the tougher subseasonal forecasting setting.
This paper complements the above prior work by developing powerful learning-based approaches that incorporate both physics-based forecast models and observational data to improve SSF over CONUS.
@@ 3 Data
Table 1 presents a description of variables used in the experiments. Climatological means of precipitation and temperature are calculated using 1971-2000 NOAA data [40]. There are many ensembles of physics-based predictions produced by forecasting systems. The North American Multi-Model Ensemble (NMME) is a collection of physics-based forecast models from various modeling centers in North America, including NOAA/NCEP, NOAA/Geophysical Fluid Dynamics Laboratory (GFDL), International Research Institute for Climate and Society (IRI), National Center for Atmospheric Research (NCAR), NASA, and Canadian Meteorological Centre [14]. NMME provides forecasts from multiple global forecast models from North American modeling centers [14]. The NMME project has two predictive periods: hindcast and forecast. A hindcast period refers to when a dynamic model re-forecasts historical events, which can help climate scientists develop and test new models to improve forecasting and evaluate model biases. In contrast, a forecast period has real-time predictions generated from dynamic models.
Table 1: Description of climate variables and their data sources. Our target climate variables for sub-seasonal forecasting are precipitation and 2-meter temperature. We use NOAA data to calculate the climatology from 1971 to 2000. We also perform linear spatial interpolation on the historical values to get values with the same support as target climate variables.
@@
tmp2m Daily average temperature at 2 meters K Globally $1^{\circ}\times 1^{\circ}$ grid 1971 to 2000 NOAA [40]
@@
Different ensemble members correspond to different initial conditions of the underlying physical model. The NCEP-CFSv2 forecasts are initialized in the following way: four initializations at times $0$, 0600, 1200, and 1800 UTC every fifth day, starting one month prior to the lead time of two weeks (Table B1 in Saha et al. [15]). NASA-GMAO is a fully coupled atmosphere–ocean–land–sea ice model, with five forecasts initialized every five days. While additional members are generated through perturbation methods closest to the beginning of each month  [^2] https://gmao.gsfc.nasa.gov/products/climateforecasts/GEOS5/DESC/init.php, not all members are initialized on different dates, meaning that the ensemble is not strictly lagged. However, NASA-GMAO members are not interchangeable, as each member is created using a distinct method.
@@ 4 Forecast tasks
The learning task can be formulated as learning a model $f_{\theta}:\mathbf{X}\to\mathbf{y}$ with parameters $\bm{\theta}$. This model $f_{\theta}$ can be a linear regression (where $\bm{\theta}$ is a set of regression weights), the mean of ensemble members (no $\bm{\theta}$ needs to be learned), a random forest (where $\bm{\theta}$ parameterizes the set of trees in the forest), a convolutional neural network (where $\bm{\theta}$ are the neural network weights), or other learned models. We consider three different forecasting tasks: regression, tercile classification, and quantile regression.
@@
$\ell_{\rm sq-err}(\theta)=\mathbb{E}[(y-f_{\theta}(\mathbf{x}))^{2}].$
@@
$\ell_{\rm CE}(\theta)=\mathbb{E}\left[\sum_{q=-1}^{1}-{{\mathbb{I}_{\{y=q\}}}}\log(f_{\theta}(\mathbf{x}))_{q}\right],$
@@ 5 Prediction methods
Our goal is to predict either the monthly average precipitation or the monthly average 2-meter temperature two weeks in advance (for example, we predict the average monthly precipitation for February on January 15). In this section, we describe the notation used for features and targets, baselines and learning methods, and how spatial features are accounted for.
@@ 5.1 Notation
We let $T$ denote the number of monthly samples and $L$ denote the number of spatial locations. We define the following variables:
@@
  • $\mathbf{z}_{l}^{(1)},\mathbf{z}_{l}^{(2)}$ represent information about longitude and latitude of location $l$, respectively; each is a vector of length $d$. More details about this representation can be found in Section 6.1.
@@
  • $y_{t,l}$ is the target – the ground truth monthly average precipitation or 2-meter temperature at the target forecast time $t+\delta_{t}$ at location $l$. For simplicity, we use a subscript $t$ for $y_{t,l}$ instead of $t+\delta_{t}$ to match with the input features notation. The same holds for the next notation.
@@
  • $\hat{s}_{m,l}$ – a 30-year climatology of a predicted climate variable, such as precipitation or temperature, at a month $m=1,\ldots,12$ and location $l$. For each location $l$ and each month $m$, it is calculated as a mean of ensemble member predictions over the training period, as defined formally in (5).
@@
    The model climatology $\hat{s}_{m,l}$ is computed using the training data. Note that we do not subtract the climatology from the input features and target variables, i.e., precipitation and temperature when training our ML models. We subtract climatology from the model outputs only when evaluating their performance, as including climatology in the inputs to our ML models improves performance. Section 6.2 provides more details on the model evaluation.
In our analyses, the number of locations $L=3274$, there are $K=24$ NCEP-CFSv2 ensemble members or $K=11$ NASA-GMAO ensemble member. The ensemble members are used as input features to the learning-based methods as they are, we do not perform any feature extraction from them. The number of observational variables is usually $P=17$. The details on these variables can be found in Section 3.
The target variable $y$ is observed from 1985 to 2020. Data from January 1985 to September 2005 are used for training (249 time steps), and data from October 2005 to December 2010 are used for validation and model selection (63 time steps). Data from 2011 to 2020 (or from 2011 to 2018 in case of NASA-GMAO data) are used for testing our methods after all model development, selection, and parameter tuning are completed.

### v4

@@ Abstract
Producing high-quality forecasts of key climate variables, such as temperature and precipitation, on subseasonal time scales has long been a gap in operational forecasting. This study explores an application of machine learning (ML) models as post-processing tools for subseasonal forecasting. Lagged numerical ensemble forecasts (i.e., an ensemble where the members have different initialization dates) and observational data, including relative humidity, pressure at sea level, and geopotential height, are incorporated into various ML methods to predict monthly average precipitation and two-meter temperature two weeks in advance for the continental United States. For regression, quantile regression, and tercile classification tasks, we consider using linear models, random forests, convolutional neural networks, and stacked models (a multi-model approach based on the prediction of the individual ML models). Unlike previous ML approaches that often use ensemble mean alone, we leverage information embedded in the ensemble forecasts to enhance prediction accuracy. Additionally, we investigate extreme event predictions that are crucial for planning and mitigation efforts. Considering ensemble members as a collection of spatial forecasts, we explore different approaches to using spatial information. Trade-offs between different approaches may be mitigated with model stacking. Our proposed models outperform standard baselines such as climatological forecasts and ensemble means. In addition, we investigate feature importance, trade-offs between using the full ensemble or only the ensemble mean, and different modes of accounting for spatial variability.
@@ 1 Introduction
The National Centers for Environmental Prediction (NCEP), part of the National Oceanic and Atmospheric Administration (NOAA), currently issues a “week 3-4 outlook” for the contiguous United States (CONUS). [^1] https://www.cpc.ncep.noaa.gov/products/predictions/WK34/ The NCEP outlooks are constructed using a combination of dynamical and statistical forecasts, with statistical forecasts based largely on how conditions in the past have varied (linearly) with indices of the El Niño-Southern Oscillation (ENSO), Madden-Julian Oscillation (MJO), and global warming (i.e., the 30-year trend). There exists great potential to advance subseasonal forecasting (SSF) using machine learning (ML) techniques. Haupt et al. [10] provides an overview of using ML methods for post-processing of numerical weather predictions. Vannitsem et al. [11] highlight the crucial role of statistical post-processing techniques, including ML methods, in national meteorological services. They discuss theoretical developments and operational applications, current challenges, and potential future directions, particularly focusing on translating research findings into operational practices. A real-time forecasting competition called the Subseasonal Climate Forecast Rodeo [12], sponsored by the Bureau of Reclamation in partnership with NOAA, USGS, and the U.S. Army Corps of Engineers, illustrated that teams using ML techniques can outperform forecasts from NOAA’s operational seasonal forecast system.
Here, we present work focused on developing ML-based forecasts that leverage lagged ensembles (i.e., an ensemble whose members are initialized from a succession of different start dates) of forecasts produced by NCEP in addition to observed data and other features. Previous studies, including successful methods in the Rodeo competition (e.g., Hwang et al. [13]), incorporate the ensemble mean as a feature in their ML systems but do not use any other information about the ensemble. In other words, variations among the ensemble members are not reflected in the training data or incorporated into the learned model. In contrast, this paper demonstrates that the full ensemble contains important information for subseasonal forecasting outside the ensemble mean. Specifically, we consider the test case of predicting monthly 2-meter temperatures and precipitation two weeks in advance over 3000 locations over the continental United States using physics-based predictions, such as NCEP-CFSv2 hindcasts [14, 15], using an ensemble of 24 distinct forecasts. We repeat this experiment for the Global Modeling and Assimilation Office from the National Aeronautics and Space Administration (NASA-GMAO) ensemble, which has 11 ensemble members [16].
@@
  • We train a variety of ML models (including neural networks, random forests, linear regression, and model stacking) that input all ensemble member predictions as features in addition to observations of geopotential heights, relative humidity, precipitation, and temperature from past months to produce new forecasts with higher accuracy than the ensemble mean; forecast accuracy is measured with a variety of metrics (Section 7). These models are considered in the context of regression, quantile regression, and tercile classification. Systematic experiments are used to characterize the influence of individual ensemble members on predictive skill (Section 8.1).
  • The collection of ML models employed allows us to consider different modes of accounting for spatial variability. ML models can account for spatial correlations among both features and targets; for example, when predicting Chicago precipitation, our models can leverage not only information about Chicago but also about neighboring regions. Specifically, we consider the following learning frameworks: (a) training a predictive model for each spatial location independently; (b) training a predictive model that inputs the spatial location as a feature and hence can be applied to any single spatial location; (c) training a predictive model for the full spatial map of temperature or precipitation – i.e., predicting an outcome for all spatial locations simultaneously. ML models present various ways to account for spatial variability, each with distinct advantages and disadvantages. Our application of model stacking (an ML technique where multiple models are combined, with their predictions used as input features for another model that produces the final prediction) allows our final learned model to exploit the advantages of each method.
@@
This paper is organized as follows: Section 2 discusses related work, Section 3 introduces data used in the experiments, Section 4 describes forecasting problems, while baselines and learning-based methods are described in Section 5, experimental setup and evaluation metrics are given in Section 6. Finally, we present our results in Section 7 and discuss them in Section 8. Conclusions and directions for future work are given in Section 9.
@@ 2 Related work
While statistical models were common for weather prediction in the early days of weather forecasting [17], forecasts using physics-based dynamic system models have been carried out since the 1980s and have been the dominant forecasting method in climate prediction centers since the 1990s [6]. Many physics-based forecast models are used both in academic research and operationally. Such systems often produce ensembles of forecasts – e.g., the result of running a physics-based simulation multiple times with different initial conditions or parameters, and are a mainstay of operational forecast centers around the globe.
@@
Some ML algorithms for subseasonal forecasting use purely observational data (i.e., not using any physics-based ensemble forecasts). He et al. [35] focuses on analyzing different ML methods, including Gradient Boosting trees and Deep Learning (DL) for SSF. They propose a careful construction of feature representations of observational data and show that ML methods are able to outperform a climatology baseline, i.e., predictions corresponding to the 30-year mean at a given location and time. This conclusion is based on comparing the relative $R^{2}$ scores for the ML approaches and climatology. Srinivasan et al. [36] proposes a Bayesian regression model that exploits spatial smoothness in the data.
Other works use the ensemble mean as a feature in their ML models. For example, in the subseasonal forecasting Rodeo [12], a prediction challenge for temperature and precipitation at weeks 3-4 and 5-6 in the western U.S. sponsored by NOAA and the U.S. Bureau of Reclamation, simple yet thoughtful statistical models consistently outperform NOAA’s dynamical systems forecasts. In particular, the winning approach uses a stacked model from two nonlinear regression models, a selection of climate variables such as temperature, precipitation, sea surface temperature, sea ice concentration, and a collection of physics-based forecast models including the ensemble mean from various modeling centers in the North American Multi-Model Ensemble (NMME). From the local linear regression with multitask feature selection model analysis, the ensemble mean is the first- or second-most important feature for forecasting, especially for precipitation. He et al. [37] perform a comparison of modern ML models that use data from the Subseasonal Experiment (SubX) project for SSF in the western contiguous United States. The experiments show that incorporating the ensemble mean as an input feature to ML models leads to a significant improvement in forecasting performance, but that work does not explore the potential value of individual ensemble members aside from the ensemble mean. Grönquist et al. [38] note that physics-based ensembles are computationally demanding to produce and propose an ML method that can input a subset of ensemble forecasts and generate an estimate of the full ensemble; they observe that the output ensemble estimate has more prediction skill than the original ensemble. Loken et al. [39] analyze the forecast skill of random forests leveraging the ensemble members for next-day severe weather prediction compared to only using the ensemble mean. However, their results only cover forecasts with a lead time of up to 48 hours, so it is unclear if their methods would have succeeded in the tougher subseasonal forecasting setting.
This paper complements the prior work above by developing powerful learning-based approaches that incorporate both physics-based forecast models and observational data to improve SSF over CONUS.
@@ 3 Data
Table 1 describes variables used in the experiments. Climatological means of precipitation and temperature are calculated using 1971-2000 NOAA data [40]. There are many ensembles of physics-based predictions produced by forecasting systems. NMME is a collection of physics-based forecast models from various modeling centers in North America, including NOAA/NCEP, NOAA/Geophysical Fluid Dynamics Laboratory (GFDL), International Research Institute for Climate and Society (IRI), National Center for Atmospheric Research (NCAR), NASA, and Canadian Meteorological Centre [14]. NMME provides forecasts from multiple global forecast models from North American modeling centers [14]. The NMME project has two predictive periods: hindcast and forecast. A hindcast period refers to when a dynamic model re-forecasts historical events, which can help climate scientists develop and test new models to improve forecasting and to evaluate model biases. In contrast, a forecast period has real-time predictions generated from dynamic models.
Table 1: Description of climate variables and their data sources. Our target climate variables for sub-seasonal forecasting are precipitation and 2-meter temperature. We use NOAA data to calculate the climatology from 1971 to 2000. We also perform linear spatial interpolation on the historical values to get values with the same resolution and support as target climate variables.
@@
Climatology tmp2m Daily average temperature at 2 meters K Globally $1^{\circ}\times 1^{\circ}$ grid 1971 to 2000 NOAA [40]
@@
Different ensemble members correspond to different initial conditions of the underlying physical model. The NCEP-CFSv2 forecasts are initialized in the following way: four initializations at times $0$, 0600, 1200, and 1800 UTC every fifth day, starting one month prior to the lead time of two weeks (Table B1 in Saha et al. [15]). NASA-GMAO is a fully coupled atmosphere–ocean–land–sea ice model, with five forecasts initialized every five days. While additional members are generated through perturbation methods closest to the beginning of each month  [^2] https://gmao.gsfc.nasa.gov/products/climateforecasts/GEOS5/DESC/init.php, not all members are initialized on different dates, meaning that the ensemble is not strictly lagged. However, NASA-GMAO members are not interchangeable, as each is created using a distinct method.
@@ 4 Forecast tasks
The learning task can be formulated as learning a model $f_{\theta}:\mathbf{X}\to\mathbf{y}$ with parameters $\bm{\theta}$. This model $f_{\theta}$ can be a linear regression (where $\bm{\theta}$ is a set of regression weights), the mean of ensemble members (no $\bm{\theta}$ needs to be learned), a random forest (where $\bm{\theta}$ parameterizes the set of trees in the forest), a convolutional neural network (where $\bm{\theta}$ is the collection of neural network weights), or other learned models. We consider three forecasting tasks: regression, tercile classification, and quantile regression.
@@
$\ell_{\rm sq-err}(\theta)=\mathbb{E}[(y-f_{\theta}(\mathbf{x}))^{2}].$ (1)
@@
$\ell_{\rm CE}(\theta)=\mathbb{E}\left[\sum_{q=-1}^{1}-{{\mathbb{I}_{\{y=q\}}}}\log(f_{\theta}(\mathbf{x}))_{q}\right],$ (2)
@@ 5 Prediction methods
Our goal is to predict either the monthly average precipitation or the monthly average 2-meter temperature two weeks in advance (for example, we predict the average monthly precipitation for February on January 15). This section describes the notation used for features and targets, baselines and learning methods, and how spatial features are accounted for.
@@ 5.1 Notation
We let $T$ denote the number of time steps used in our analysis, and $L$ denote the number of spatial locations. We define the following variables:
@@
  • $\mathbf{z}_{l}^{(1)},\mathbf{z}_{l}^{(2)}$ represent information about longitude and latitude of location $l$, respectively; each is a vector of length $d$. More details about this representation, called positional encoding (PE), can be found in Section 6.1.
@@
  • $y_{t,l}$ is the target – the ground truth monthly average precipitation or 2-meter temperature at the target forecast time $t+\delta_{t}$ at location $l$, where $\delta_{t}=14$ days is our forecast horizon. For simplicity, we use a subscript $t$ for $y_{t,l}$ instead of $t+\delta_{t}$ to match with the input features notation. The same holds for our ensuing definitions.
@@
  • $\hat{s}_{m,l}$ – a 30-year climatology of a predicted climate variable, such as precipitation or temperature, at a month $m=1,\ldots,12$ and location $l$. For each location $l$ and each month $m$, it is calculated as a mean of ensemble member predictions over the training period, as defined formally in Eq. 7.
@@
    The model climatology $\hat{s}_{m,l}$ is computed using the training data. Note that we do not subtract the climatology from the input features and target variables, i.e., precipitation and temperature, when training our ML models. We subtract climatology from the model outputs only when evaluating their performance, as including climatology in the inputs to our ML models during training improves performance. Section 6.3 provides more details on the model evaluation.
In our analyses, the number of locations is $L=3274$, there are $K=24$ NCEP-CFSv2 ensemble members or $K=11$ NASA-GMAO ensemble members. The ensemble members are used as input features to the learning-based methods as they are, we do not perform any feature extraction from them. The number of observational variables is usually $P=17$. The details on these variables can be found in Section 3 and Section 6.2.
The target variable $y$ is observed from 1985 to 2020. Data from January 1985 to September 2005 are used for training (249 time steps), and data from October 2005 to December 2010 are used for validation and model selection (63 time steps). Data from 2011 to 2020 (or from 2011 to 2018 in the case of NASA-GMAO data) are used to test our methods after all model development, selection, and parameter tuning are completed.